OpenStudy (anonymous):

The state pension increased by 2% to £78.03 What was it before the increase show your working

OpenStudy (vishweshshrimali5):

Let old income = x then \[x + \frac{2}{100} * x = 78.03\] => \[\frac{102x}{100} = 78.03\] => \[102x = 7803\] => \[x = \frac{7803}{102}\] This is the original one.

OpenStudy (anonymous):

thnx for answering but can u just tell me what the price was before the increase as i find this method confusing so if i know the answer the ill be able to make sense maybe

OpenStudy (anonymous):

i got it